describe how to prepare exactly 250.0 mL of 0.125 M of BaCl2 ∙ 2H2O (this is a hydrated compound; the mass of the water must be included in the molar mass of the compound but the water of hydration does not affect the ions formed when the solid is dissolved)
1
Expert's answer
2019-09-24T05:57:46-0400

Molar mass of "BaCl_2•2H_2O=244.26 \ g\ mol^{-1}"

"Molarity=\frac{mass}{molar\ mass\times Volume(in\ l)}"

"mass=molarity\times molar\ mass\times volume\\ \ \ \ \ =0.125 \times244.26\times0.250=7.633\ g"

So "7.633 \ g" of this salt is added into water so that final volume becomes 250 ml